# Video Prediction Policy: A Generalist Robot Policy with Predictive Visual Representations

### v1

@@ 3.3 Baseline Implementations
Table 9: We outline the dataset scales and sample ratios used for training our manipulation text-guided video prediction model. Following [22], we exclude 5,558 bridge trajectories and 2,048 something-something-v2 trajectories during training, reserving them for validation. For all other datasets, 3% of the trajectories are excluded and used as validation datasets.
(a) moving bottle away from bottle
(b) moving silicone towards the camera
(c) moving toy closer to plastic glass
(d) pouring liquid into a cup
(e) pouring milk into glass
(f) turning the camera right while filming jeep
(g) pretending to pick a tennis ball up
(h) pulling toy car from left to right
(i) putting something similar
(j) tearing paper into two pieces
(a) open bottom drawer
(b) place apple into bottom drawer
(c) place eraser in purple bowl
(d) place the white sponge in the ceramic cup
(e) put bowl on plate cardboard fence
(f) put eggplant into pot or pan
(g) Place the grasped object into the drawer
(h) close the drawer
(i) Put the grey circle on the red pole
(j) open the drawer
(k) pick carrot
(l) pick the blue block
(m) Put the green block above the red block
(n) Place the purple onion onto the basket
(o) pick the yellow ball to the empty hole of the board
(p) Receive the orange from hand and place it into the basket
(a) place the grasped object in the drawer.
(b) close the drawer.
(c) pick the mango and place in blue plate
(d) stack the green block on red block

### v2

@@ B.1 Datasets Sample Ratios
Table 8: We outline the dataset scales and sample ratios used for training our manipulation text-guided video prediction model. Following (Gu et al., 2023), we exclude 5,558 bridge trajectories and 2,048 something-something-v2 trajectories during training, reserving them for validation. For all other datasets, 3% of the trajectories are excluded and used as validation datasets.
B.2 Quantitative result on Prediction Quality
Our TVP models successfully predict future frames on validation datasets across diverse manipulation tasks, with some prediction results visualized in Appendix B.3. Additionally, we evaluate the quantitative FVD metric (Unterthiner et al., 2018) on the bridge datasets (Ebert et al., 2021), following the evaluation settings in Seer (Gu et al., 2023). The results are shown in Table 9. Surprisingly, our model easily outperforms the previous TVP model. We attribute this improvement to our use of the pre-trained video foundation model SVD (Blattmann et al., 2023a), which the earlier TVP model did not leverage, giving us a significant advantage.
Bridge VideoFusion Tune-A-Video Seer VPP
[…]
Table 9: Quantitative evaluation of prediction quality on bridge datasets. The results of VideoFusion (Luo et al., 2023), Tune-A-Video (Wu et al., 2023b), Seer (Gu et al., 2023) are copied from (Gu et al., 2023).
B.3 More Visualization of Complete Prediction Results
We present additional visualizations of prediction results from our fine-tuned manipulation TVP model. Predictions on human manipulation datasets are displayed in Figure 8, and those on robotic manipulation datasets are illustrated in Figure 10. All trajectories are sampled from the validation datasets and are predicted using the same manipulation TVP model. Each sample was denoised in 30 steps using classifier-free guidance set at 7.5, as described in (Gu et al., 2023). Our TVP model predicts a horizon of 16, and we visualize 8 frames at a skip step of 2 due to space constraints.
B.4 More Visualizations of Predictive Representations
Appendix C More Details for Experiments
C.1 Structure details
We provide the VPP architecture and hyperparameter setting details in four evaluate environments, as shown in Table 13. The transformer block in TVP follows the setting in (Blattmann et al., 2023a), and the rest of the hyperparameter in Diffusion Transformer follows the work (Reuss et al., 2024).
C.2 More ablation
[…]
Ablation 1 on the video former entails the removal of the Temporal-attn module from the Video Former while maintaining all other configurations same as VPP. The results, displayed in Table 12, demonstrate that the Temporal-attn module could enhance the temporal comprehension capabilities of the Video Former.
Ablation 2 on the number of denoising steps introduces a 2-step denoising process in the TVP to derive the predictive visual representation. The outcomes are summarized in Table 12, revealing that the 2-step process did not yield superior performance. We hypothesize this is because a single denoising step suffices to generate an effective representation for trajectory prediction in our configuration. Additionally, the 2-step denoising process nearly doubles the inference time and reduces the control frequency by half. Due to these factors, we opted for a one-step direct encoder in our main experiments.
[…]
Ablations on using different layers of features The average task completion length are listed in Table 10.
Ablations on using different diffusion time-step The average task completion length are listed in Table 11.
Calvin abc-d Layer-3 Layer-6 Layer-9 Layer-12 VPP
Avg. Len 3.72 3.88 4.29 4.05 4.33
Table 10: Ablations on different layers of features.
Calvin abc-d Time-step 10 Time-step 20 Time-step 30
Avg. Len 4.21 4.33 4.25
Table 11: Ablations on the Use of different diffusion time-step.
Method Tasks completed in a row
1 2 3 4 5 Avg. Len $\uparrow$
VPP(Ours) 0.965 0.909 0.866 0.820 0.769 4.33
VPP(Single-view) 0.909 0.815 0.713 0.620 0.518 3.58
Ablation.1 0.949 0.900 0.839 0.780 0.714 4.18
Ablation.2 0.951 0.904 0.840 0.777 0.718 4.19
Table 12: More ablation studies.
C.3 Baseline Implementations
The baseline methods, including RT-1 (Brohan et al., 2022), GR-1 (Wu et al., 2023a), and Diffusion Policy (Chi et al., 2023), are implemented based on their official repositories. For comparison with Susie (Black et al., 2023) in both the Metaworld and real-world manipulation scenarios, we adopt InstructPix2Pix (Brooks et al., 2023) as the future frame predictor and use an image-goal Diffusion Policy (Chi et al., 2023) to generate the state sequence.